# Online Fair Allocations with Binary Valuations and Beyond

### v1

@@ Abstract
In an online fair allocation problem, a sequence of indivisible items arrives online and needs to be allocated to offline agents immediately and irrevocably. In our paper, we study the online allocation of either goods or chores. We adapt popular fairness notions, including envy-freeness up to one item (EF1) and maximin share fairness (MMS), and use utilitarian social welfare (USW) to measure efficiency. For both settings of items, we present a series of positive results regarding the existence of fair and efficient allocations with widely studied classes of binary and bivalued valuation/cost functions. Additionally, we complement our results by constructing some counterexamples to establish our results as among the best guarantees possible.
@@ 1 Introduction
How to allocate a set of indivisible items fairly among a set of agents is a hot topic in various research communities, including algorithmic game theory and artificial intelligence. For decades, there have been tremendous ongoing efforts to design efficient algorithms for finding fair allocations regarding envy-free up to one item (EF1), maximin share fairness (MMS), and some other fairness; detailed progress and results can be found in [1]. Most prior work focuses on offline setting, where the set of items is known upfront. In many practical situations, decisions about the allocation of items must be made online without or with limited knowledge about the future. A paradigmatic example is the food banks problem [2]. In this scenario, food comes online without prior knowledge of future food, and the food is perishable; it must be allocated upon arrival to one of the agents.
In this paper, we study the online problem of allocating a set of items that arrive one by one to a set of agents that are known in advance. When an item arrives, its values to the agents are revealed, and an immediate and irrevocable decision must be made about which agent to assign the item. There has been a line of work about online fair allocation, where they consider either additive binary valuations [3] or normalized valuations [4, 5, 6]. Few non-trivial results about the approximation fairness guarantee of online goods allocation beyond additive binary valuations and online chores allocation are known when there is no prior information about the items. In our paper, we would like to fill this gap and consider the online allocation model without any information about future items, where we discuss the goods and chores settings, respectively. Our primary research question is
Can we design online algorithms that are fair and efficient, for general valuation functions beyond additive binary, in the online allocation problem?
@@ 1.1 Our Contributions
Valuations $N$ Fairness/ Efficiency LB UB
@@
[…]
USW $\times$ (Prop 2)
(a) Goods setting
Costs $N$ Fairness/ Efficiency LB UB
[…]
(b) Chores setting
Table 1: A summary of our main results. "$\times$" indicates inapproximability, $N$ indicates the number of agents. "LB" indicates lower bound, and "UB" indicates upper bound.
There are $n$ agents, and each indivisible item comes online. We study the goods and chores settings, and introduce non-wastefulness (NW) and completeness constraints to the goods and chores settings, respectively. For different classes of valuation or cost functions, including submodular/supermodular binary, additive bi-valued, and so on, our main results are summarized in Table 1.
In the goods setting, we first show the inapproximability of EF1 and MMS with general additive valuations, even for only two agents with additive tri-valued valuations. This indicates that we can only focus our efforts on restricted valuation functions.
  • For submodular binary valuation functions, we propose the Marginal-Greedy Algorithm (Algorithm 2) to compute allocations that simultaneously satisfy NW, $\frac{1}{2}$-EF1, $\frac{1}{2}$-MMS, and $\frac{1}{2}$-max-USW. Further, we show that no deterministic algorithm can guarantee an NW allocation that is $\alpha$-EF1, $\alpha$-MMS, or $\alpha$-max-USW for any $\alpha>\frac{1}{2}$.
  • For two agents with additive bi-valued valuation functions, we design the Adapted Envy-Graph Procedures (Algorithm 3) to compute NW, $\frac{1}{2}$-EF1 and $\frac{1}{3}$-MMS allocations. Further, we show that no deterministic algorithm can guarantee a NW allocation that is $\alpha$-EF1 for any $\alpha>\frac{1}{2}$ or $\beta$-MMS for any $\beta>\frac{1}{3}$, and no deterministic algorithm can achieve NW, $\alpha$-EF1 or $\alpha$-MMS, and $\gamma$-max-USW for any $\alpha,\gamma>0$.
  • When one agent has the additive bi-valued valuation function and the remaining agents have identical additive binary valuation functions, we design the Adapted-Picking Algorithm (Algorithm 6) to compute an allocation that simultaneously satisfies NW, EF1, and MMS. Further, we show that no deterministic algorithm can compute an NW allocation that is EF1 or MMS guarantee $\alpha$-max-USW for any $\alpha>0$.
In the chores setting, we first demonstrate the inapproximability of EF1 for additive tri-valued cost functions, which hinders our ability to explore more general cost functions. Next, we investigate the restricted cost functions, such as additive binary cost functions.
  • For additive binary cost functions, we propose the Compelled-Greedy Algorithm (Algorithm 7) to compute a complete allocation that simultaneously satisfies EF1, MMS, and min-USC.
  • For supermodular binary cost functions, we prove that no deterministic algorithm can compute a complete allocation that is $\alpha$-EF1, $\alpha$-MMS, or $\alpha$-min-USC for any $\alpha\geq 1$.
  • When two agents have additive bi-valued cost functions, we design the Adapted Chores Envy-Graph Procedures (Algorithm 8) to compute an allocation that satisfies completeness, $2$-EF1, and $\frac{5}{3}$-MMS. Further, we show that no deterministic algorithm can guarantee a complete allocation that is $\alpha$-EF1 for any $\alpha<2$ or $\beta$-MMS for any $\beta<\frac{3}{2}$, and no deterministic algorithm can guarantee a complete allocation that is $\alpha$-EF1 or $\beta$-MMS, and $\gamma$-min-USC for any $\alpha,\gamma\geq 1$ and $\beta<2$.
  • When one agent has the additive bi-valued cost function and the remaining agents have additive binary cost functions, we design the Adapted-Chores-Picking Algorithm (Algorithm 11) to compute an allocation that simultaneously satisfies completeness, EF1 and MMS. Further, no deterministic algorithm can compute a complete allocation that is $\alpha$-EF1 or MMS guarantee $\beta$-min-USC for any $\alpha,\beta\geq 1$.
@@ 1.2 Related Work
In the majority of the literature, there are two kinds of online arrivals: The former kind is that agents arrive over time [7, 8, 9, 10], and the latter is that items arrive over time, like [3], which is also the topic of our paper. Our work is mainly related to the study of [3] and [11], where items arrive one by one without any future information known in advance. Aleksandrov et al. [3] consider the food bank problem when all agents only have the additive binary valuation functions and design two simple mechanisms to guarantee ex-ante envy-freeness. In our paper, we study beyond additive binary valuation functions in the goods setting. Hosseini et al. [11] study the class fairness in online matching for indivisible and divisible items, and present a series of positive results. The model of indivisible items in their paper can be seen as a special case of our model with submodular binary valuation functions in our paper, which will be discussed later. In other words, we study a more general setting and derive the approximation results that strengthen their results.
Some work studies about online fair allocation with some prior information of future items, like the normalized valuations [5, 6], the rank of a new arrival item [12], and other information [13, 14, 15]. Zhou et al. [5] study the online allocation of indivisible goods or chores with normalized valuations and design several algorithms only to guarantee the final allocation that approximates maximin share, which is in contrast to our setting, where our online algorithm ensures the approximation ratio in any round. Benade et al. [6] study the trade-off between fairness and efficiency in allocating the indivisible items under different types of adversaries with normalized valuations. Benade et al. [12] study the allocation of indivisible items with partial information, i.e., the rank of a new arrival item with respect to the past items under different assumptions of agents’ valuation distributions. He et al. [13] study the minimum adjustments (the reallocation of previously allocated items) to ensure envy-freeness up to one item for online indivisible items allocation with or without the information where the values of future items are observed. Elkind et al. [14] and Cookson et al.[15] study the temporal allocation of online indivisible items where the arriving information of items, like values of items, is known and give several positive results. With some known information, we also explore online fair allocation, and the relevant results are presented in Appendix B.
@@
When agents have additive valuation functions, there is a rich body of literature that discusses EF1 or MMS. Please see the survey of [1] for an overview. Here, we list some papers most relevant to our work about offline fair division with submodular binary valuation functions or supermodular binary cost functions. When agents have submodular binary valuations in the goods setting, Benabbou et al. [16] study the compatibility of fairness and efficiency and show that a maximizing utilitarian social welfare allocation that is EF1 can be efficiently computed, and Barman et al. [17] consider the fairness notion - MMS and prove that a maximizing utilitarian social welfare allocation that satisfies MMS can be computed in polynomial time. When agents have supermodular binary cost functions in the chores setting, Barman et al. [18] show that an EF1 or MMS allocation that minimizes the social cost can be computed in polynomial time.
@@ 2 Preliminaries
For each natural number $s\in\mathbb{N}$, let $[s]=\{1,\ldots,s\}$. Let $T=\{e_{1},\ldots,e_{t}\}$ be the set of indivisible items arriving one by one and $N=\{1,\ldots,n\}$ be the set of offline agents. When an online item $e_{t}$ arrives, its incident values or costs to agents are revealed. When items are goods, each agent is endowed with a valuation function $v_{i}:2^{T}\rightarrow\mathbb{R}_{\geq 0}$. When items are chores, each agent is endowed with a cost function $c_{i}:2^{T}\rightarrow\mathbb{R}_{\geq 0}$. We assume that $v_{i}(\emptyset)=0$ ($c_{i}(\emptyset)=0$), and monotone, $v_{i}(X)\leq v_{i}(Y)$ ($c_{i}(X)\leq c_{i}(Y)$) for all $X\subseteq Y\subseteq M$. For any item $e\in T$ and any subset $X\subseteq T$, we denote by $\Delta^{i}_{X}(e)=v_{i}(X\cup\{e\})-v_{i}(X)$ or $c_{i}(X\cup\{e\})-c_{i}(X)$ the marginal value or cost of an item $e\in M$ in the set $X$. $v_{i}$ is additive if the values of any set of items $S$ is the sum of the value of each item in $S$, that is, $v_{i}(X)=\sum_{e\in X}v_{i}(e)$. $v_{i}$ is submodular if $v_{i}(X)+v_{i}(Y)\geq v_{i}(X\cup Y)+v_{i}(X\cap Y)$ for any two sets $X,Y\subseteq T$. $v_{i}$ is supermodular if $v_{i}(X)+v_{i}(Y)\leq v_{i}(X\cup Y)+v_{i}(X\cap Y)$ for any two sets $X,Y\subseteq T$. There are similar arguments for $c_{i}$, and we omit them. Let $T^{k}$ be the set of items that have arrived by round $k\in[t]$. We denote an allocation by $A^{k}=(A_{1}^{k},\ldots,A_{n}^{k})$, where $A_{i}^{k}\subseteq T^{k}$ is the bundle of items that are allocated to agent $i$, and $A_{i}^{k}\cap A_{j}^{k}=\emptyset$ for any $i,j\in N$. Collectively, $(T,N,(v_{i})_{i\in N})$ forms an instance of online goods fair allocation and $(T,N,(c_{i})_{i\in N})$ forms an instance of online chores fair allocation. For a singleton set $\{e\}$, we will use $v_{i}(e)$ ($c_{i}(e)$) as a shorthand for $v_{i}(\{e\})$ ($c_{i}(\{e\})$).
Deterministic online algorithms
When one item arrives, the values (costs) of this item for all agents are revealed. The online algorithm must make an immediate and irrevocable decision to allocate the item to one agent or throw it away deterministically.
Our goal is to design the deterministic algorithms to compute allocations guaranteeing the desired approximation fairness or efficiency (approximate EF1, MMS, USW, USC) (Definitions 1, 2, 3, 8, 9, 10) and constraint (non-wastefulness for goods (Definition 4) and completeness for chores (Definition 7)) at the end of each round. In contrast, our impossibility results will hold even if the desired guarantees are required to hold only at the end. Let us demonstrate the deterministic online allocation procedure with the following example.
$t=1$ $t=2$ $t=3$ $t=4$ $t=5$
agent $1$ $6$ $4$ $5$ 9 3
agent 2 $5$ $8$ $10$ 2 5
agent 3 $12$ 6 6 8 4
@@
Let us consider the following example to demonstrate the deterministic online allocation procedure. Suppose that there are three agents who have additive valuation functions, and the arriving items are goods. The value of each item is shown in Table 2. Now we have a deterministic online algorithm named $\mathsf{OALG}$. In round $t=1$, $\mathsf{OALG}$ allocates $e_{1}$ to agent 1 after the values of this item are revealed. For allocation $A^{1}$, it is EF1, MMS and $\frac{1}{2}$-max-USW. In round $t=2$, by $\mathsf{OALG}$, $e_{2}$ is allocated to agent 2, and allocation $A^{2}$ is EF1, MMS, and $\frac{7}{10}$-max-USW. In round $t=3$, by $\mathsf{OALG}$, agent 3 picks item $e_{3}$ and allocation $A^{3}$ is EF1, MMS, and $\frac{2}{3}$-max-USW. In round $t=4$, by $\mathsf{OALG}$, agent 1 picks item $e_{4}$ and allocation $A^{4}$ is $\frac{3}{4}$-EF1, $\frac{3}{4}$-MMS, and $\frac{29}{39}$-max-USW. In round $t=5$, by $\mathsf{OALG}$, agent 1 picks item $e_{5}$ and allocation $A^{5}$ is $\frac{1}{2}$-EF1, $\frac{1}{2}$-MMS, and $\frac{8}{11}$-max-USW. Then, no further item will arrive and we say that $\mathsf{OALG}$ achieves $\min\{\frac{1}{2},\frac{3}{4},1\}=\frac{1}{2}$-EF1, $\min\{\frac{1}{2},\frac{3}{4},1\}=\frac{1}{2}$-MMS, and $\min\{\frac{1}{2},\frac{2}{3},\frac{7}{10},\frac{8}{11},\frac{29}{39}\}=\frac{1}{2}$-max-USW.

### v2

# Yuanyuan Wang1    Tianze Wei2

@@ Abstract
In an online fair allocation problem, a sequence of indivisible items arrives online and needs to be allocated to offline agents immediately and irrevocably. In our paper, we study the online allocation of either goods or chores. We employ popular fairness notions, including envy-freeness up to one item (EF1) and maximin share fairness (MMS) to capture fairness, and utilitarian social welfare (USW) to measure efficiency. For both settings of items, we present a series of positive results regarding the existence of fair and efficient allocations with widely studied classes of additive binary and personalized bi-valued valuation/cost functions. Furthermore, we complement our results by constructing counterexamples to establish our results as among the best guarantees possible.
@@ 1 Introduction
The fair division of indivisible items is a prominent topic in algorithmic game theory and artificial intelligence, with practical applications Budish (2011); Goldman and Procaccia (2015). Most prior work has focused on offline settings, where the information of all items is known a priori. In practice, items are often online, and waiting for all resources to appear can significantly impede progress. We must proactively allocate resources in a timely manner for better outcomes. A paradigmatic example is the food banks problem Lee et al. (2019), where food arrives online without prior knowledge of future items. Since the food is perishable, it must be allocated upon arrival to some agent. Motivated by these real-world scenarios, we define and discuss the following problem:
(Online Fair Division) There are $n$ agents that are fixed and known in advance. The items arrive online (the number of items is unknown). Upon the arrival of an item, its values to all agents are revealed, and we must allocate it to some agent or discard it immediately and irrevocably. At all times, the allocation among agents must be fair.
If the online algorithm knows additional information about all future items beforehand, it somewhat reduces the uncertainty of the online model, allowing for more effective decision-making and improved outcomes. Zhou et al. (2023); Benadè et al. (2024) study the normalized valuations, i.e., online algorithms know the total value of future items. Benadè et al. (2022) studies the online fair allocation, foreseeing the rank of a new arrival item with respect to the past items under different assumptions of agents’ valuation distributions. And there are other studies about additional information He et al. (2019); Elkind et al. (2024); Cookson et al. (2024). For an independent interest, we also examine the online setting with additional information, with relevant results presented in the Appendix.
In numerous practical applications, such as vaccine allocation Rey et al. (2023), the online algorithm operates without prior knowledge of future items. Aleksandrov et al. (2015) studies the online fair division with additive binary valuations and shows that an envy-free up to one item (EF1) allocation exists. However, there are a few non-trivial approximation fairness results of online goods allocation beyond additive binary valuations and online chores allocation. In our paper, we aim to address this gap and study the online allocation model without any information about future items. We discuss the goods and chores settings, respectively. Our primary research question is:
Can we design deterministic online algorithms that are fair and efficient, for valuation functions beyond additive binary, in the online allocation problem?
@@ 1.1 Our Contributions
We study two settings: goods and chores, in the online fair indivisible items allocation model.
In the goods setting, for different classes of valuations, including binary and additive personalized bi-valued, our main results are summarized in Table 1. We introduce the non-wastefulness (NW) notion in this setting. First, we show the inapproximability of EF1 and MMS with general additive valuations, even for only two agents with additive personal tri-valued valuations. Then, we study some restricted valuation functions, such as submodular binary valuation functions.
Valuations
Number
of agents
Fairness/
Efficiency
LB UB
Additive
(personalized tri-valued)
[…]
Submodular
binary
$n$ EF1
$\frac{1}{2}$
(Thm 2)
$\frac{1}{2}$
(Thm 3)
[…]
MMS
$\frac{1}{2}$
(Thm 2)
$\frac{1}{2}$
(Thm 3)
[…]
USW
$\frac{1}{2}$
(Thm 2)
$\frac{1}{2}$
(Thm 3)
[…]
Additive personalized
bi-valued
22 EF1
$\frac{1}{2}$
(Thm 4)
$\frac{1}{2}$
(Thm 5)
[…]
MMS
$\frac{1}{3}$
(Thm 4)
$\frac{1}{3}$
(Thm 5)
[…]
Identical additive
binary
+ personalized bi-valued
[…]
USW $\times$ (Prop 3)
Table 1: Goods setting. LB: lower bound. UB: upper bound. $\times"$ indicates inapproximability.
  • For submodular binary valuation functions, we propose the Marginal-Greedy Algorithm (Algorithm 2) to compute an allocation that simultaneously satisfies NW, $\frac{1}{2}$-EF1, $\frac{1}{2}$-MMS, and $\frac{1}{2}$-max-USW, and we show that these approximation ratios are all tight.
  • For two agents with additive personalized bi-valued valuation functions, we design the Adapted Envy-Graph Procedure (Algorithm 3) to compute an NW allocation satisfying $\frac{1}{2}$-EF1 and $\frac{1}{3}$-MMS simultaneously. Further, we show that no deterministic algorithm can guarantee an NW allocation that is $\alpha$-EF1 or $\beta$-MMS for any $\alpha>\frac{1}{2}$, $\beta>\frac{1}{3}$, and no deterministic algorithm can compute an NW allocation where $\alpha$-EF1 (or $\alpha$-MMS) is compatible with $\beta$-max-USW for any $\alpha,\beta>0$.
  • When one agent has the additive personalized bi-valued valuation function and the remaining agents have identical additive binary valuation functions, we design the Adapted-Picking Algorithm (Algorithm 7) to compute an allocation satisfying NW, EF1, and MMS. Further, we show that no deterministic algorithm can compute an NW allocation that is EF1 (or MMS) and $\alpha$-max-USW for any $\alpha>0$.
Costs
Number
of agents
Fairness/
Efficiency
LB UB
Additive
(personalized tri-valued)
[…]
Additive
binary
@@ 22 EF1
Additive personalized
bi-valued
22 EF1
22
(Thm 11)
22
(Thm 10)
[…]
MMS
$\frac{3}{2}$
(Thm 11)
$\frac{5}{3}$
(Thm 10)
[…]
Additive binary
+ personalized bi-valued
@@
Table 2: Chores setting. LB: lower bound. UB: upper bound. $\times"$ indicates inapproximability.
In the chores setting, we introduce the completeness constraints. We first show the inapproximability of EF1 even for two agents with additive personalized tri-valued cost functions, which blocks our way of studying more general cost functions. Next, we investigate the restricted cost functions, such as additive binary cost functions.
  • For additive binary cost functions, we propose the Compelled-Greedy Algorithm (Algorithm 8) to compute a complete allocation that simultaneously satisfies EF1, MMS, and min-USC. For supermodular binary cost functions, we prove that no deterministic algorithm can compute a complete allocation that is $\alpha$-EF1 (or $\alpha$-MMS or $\alpha$-min-USC) for any $\alpha\geq 1$.
  • When two agents have additive personalized bi-valued cost functions, we design the Adapted Chores Envy-Graph Procedure (Algorithm 9) to compute an allocation that satisfies completeness, $2$-EF1, and $\frac{5}{3}$-MMS. Further, we show that no deterministic algorithm can guarantee a complete allocation that is $\alpha$-EF1 for any $\alpha<2$ or $\beta$-MMS for any $\beta<\frac{3}{2}$, and no deterministic algorithm can guarantee a complete allocation that is $\alpha$-EF1 (or $\beta$-MMS) and $\gamma$-min-USC for any $\alpha,\gamma\geq 1$ and $\beta<2$.
  • When one agent has the additive personalized bi-valued cost function and the remaining agents have additive binary cost functions, we design the Adapted-Chores-Picking Algorithm (Algorithm 13) to compute an allocation satisfying completeness, EF1, and MMS. Further, no deterministic algorithm can compute a complete allocation that is $\alpha$-EF1 (or MMS) and $\beta$-min-USC for any $\alpha,\beta\geq 1$.
@@ 1.2 Related Work
In the majority of the literature, there are two kinds of online arrivals: The former kind is that agents arrive over time Kash et al. (2014); Friedman et al. (2015, 2017); Li et al. (2018), and the latter is that items arrive over time, like Aleksandrov et al. (2015), which is also the topic of our paper. Our work is mainly related to the study of Aleksandrov et al. (2015) and Hosseini et al. (2024), where items arrive one by one without any future information known in advance. Aleksandrov et al. (2015) consider the food bank problem when all agents only have the additive binary valuation functions and design two simple mechanisms to guarantee ex-ante envy-freeness. In our paper, we study beyond additive binary valuation functions in the goods setting. Hosseini et al. (2024) study the class fairness in online matching for indivisible and divisible items, and present a series of positive results. The model of indivisible items in their paper can be seen as a special case of our model with submodular binary valuation functions in our paper, which will be discussed later. In other words, we study a more general setting and derive the approximation results that strengthen their results.
@@
When agents have additive valuation functions, there is a rich body of literature that discusses EF1 or MMS. Please refer to the survey by Amanatidis et al. (2023) for an overview. Here, we highlight some papers most relevant to our work on offline fair division with submodular binary valuation functions or supermodular binary cost functions. When agents have submodular binary valuations in the goods setting, Benabbou et al. (2021) study the compatibility of fairness and efficiency and show that a maximum utilitarian social welfare allocation that is EF1 can be efficiently computed, and Barman and Verma (2021) consider the fairness notion - MMS and prove that a maximum utilitarian social welfare allocation that satisfies MMS can be computed in polynomial time. When agents have supermodular binary cost functions in the chores setting, Barman et al. (2023) show that an EF1 or MMS allocation minimizing the social cost can be computed in polynomial time.
@@ 2 Preliminaries
For each natural number $s\in\mathbb{N}$, let $s]=\{1,\ldots,s\}$. Let $T=\{e_{1},\ldots,e_{t}\}$ be the set of indivisible items arriving one by one and $N=\{1,\ldots,n\}$ be the set of offline agents. When an online item $e_{t}$ arrives, its values or costs to all agents are revealed. When items are goods, each agent is endowed with a valuation function $v_{i}:2^{T}\rightarrow\mathbb{R}_{\geq 0}$. When items are chores, each agent is endowed with a cost function $c_{i}:2^{T}\rightarrow\mathbb{R}_{\geq 0}$. We assume that $v_{i}(\emptyset)=0$ ($c_{i}(\emptyset)=0$), and monotone, $v_{i}(X)\leq v_{i}(Y)$ ($c_{i}(X)\leq c_{i}(Y)$) for all $X\subseteq Y\subseteq T$. For any item $e\in T$ and any subset $X\subseteq T$, we denote by $\Delta^{i}_{X}(e)=v_{i}(X\cup\{e\})-v_{i}(X)$ or $c_{i}(X\cup\{e\})-c_{i}(X)$ the marginal value or cost of an item $e\in T$ in the set $X$. $v_{i}$ is additive if the value of any set of items $S$ is the sum of the value of each item in $S$, that is, $v_{i}(X)=\sum_{e\in X}v_{i}(e)$. $v_{i}$ is submodular if $v_{i}(X)+v_{i}(Y)\geq v_{i}(X\cup Y)+v_{i}(X\cap Y)$ for any two sets $X,Y\subseteq T$. $v_{i}$ is supermodular if $v_{i}(X)+v_{i}(Y)\leq v_{i}(X\cup Y)+v_{i}(X\cap Y)$ for any two sets $X,Y\subseteq T$. There are similar arguments for $c_{i}$, and we omit it. Let $T^{k}$ be the set of items that have arrived by round $k\in[t]$. We denote an allocation by ${\bf A}^{k}=(A_{1}^{k},\ldots,A_{n}^{k})$, where $A_{i}^{k}\subseteq T^{k}$ is the bundle of items that are allocated to agent $i$, and $A_{i}^{k}\cap A_{j}^{k}=\emptyset$ for any $i,j\in N$. Let $\Pi_{n}(T^{k})$ be the set of all allocations of $T^{k}$. Collectively, $(T,N,(v_{i})_{i\in N})$ forms an instance of online goods fair allocation and $(T,N,(c_{i})_{i\in N})$ forms an instance of online chores fair allocation. For a singleton set $\{e\}$, we will use $v_{i}(e)$ ($c_{i}(e)$) as a shorthand for $v_{i}(\{e\})$ ($c_{i}(\{e\})$).
Deterministic online algorithms.
In the online setting, the items in $T$ arrive one by one in an arbitrary order, where $|T|$ is unknown. When an item arrives, the values (costs) of this item for all agents are revealed. The online algorithm must make an immediate and irrevocable decision to allocate the item to one agent or discard it deterministically.
Our objective is to design deterministic algorithms to compute allocations guaranteeing the desired approximation fairness or efficiency, including EF1, MMS, and USW/USC (Definitions 1, 2, 3, 8, 9, and 10), as well as constraints, including non-wastefulness (Definition 4) and completeness (Definition 7), at the end of each round. In contrast, our impossibility results will hold even if the desired guarantees are required to hold only at the end. For a better understanding of our online model, let us demonstrate the deterministic online allocation procedure through the following example.
$e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$
agent 11 6 88 10 66
agent 2 13 8 5 0
@@
Suppose that there are two agents who have additive valuation functions, and the arriving items are goods. The value of each item is shown in Table 3, and note that the value of each item is only revealed after it arrives. Now we have a deterministic online algorithm named $\mathsf{OALG}$. In round $t=1$, $\mathsf{OALG}$ allocates $e_{1}$ to agent 1, and the allocation ${\bf A}^{1}$ is EF1, MMS, and $\frac{6}{13}$-max-USW. In round $t=2$, $\mathsf{OALG}$ allocates $e_{2}$ to agent 2, and the allocation ${\bf A}^{2}$ is EF1, MMS, and $\frac{2}{3}$-max-USW. In round $t=3$, $\mathsf{OALG}$ allocate $e_{3}$ to agent 2, and the allocation $\textbf{A}^{3}$ is $\frac{3}{4}$-EF1, $\frac{3}{5}$-MMS, and $\frac{19}{31}$-max-USW. In round $t=4$, $\mathsf{OALG}$ discards $e_{4}$, and the allocation $\textbf{A}^{4}$ is $\frac{3}{4}$-EF1, $\frac{3}{7}$-MMS, and $\frac{19}{37}$-max-USW. Then, no further item arrives and we say that $\mathsf{OALG}$ achieves $\min\{\frac{3}{4},1\}=\frac{3}{4}$-EF1, $\min\{\frac{3}{7},\frac{3}{5},1\}=\frac{3}{7}$-MMS, and $\min\{\frac{6}{13},\frac{19}{37},\frac{19}{31},\frac{2}{3}\}$$\frac{6}{13}-$max-USW.